OpenStudy (anonymous):

Consider rectangle BCDE, if m<4=35, find the measure of each numbered angles.

OpenStudy (anonymous):



OpenStudy (anonymous):

m<3= 90-m<4 m<3=m<1 m<2= 180-(m<3+m<1)